[Date Prev][Date Next][Thread Prev][Thread Next][Date Index][Thread Index]

Mars Pathfinder Update - September 11, 1997

11 September 1997, 7:30 p.m. PDT

We have completed another day of operations at the Carl
Sagan Memorial Station. Sunrise occurred about 9:30 PDT
and sunset will occur about 9:30 PM this evening.

Last night we received 10 megabits of science and rover
data which included meteorological data from the ASI/MET
instrument, wind sock instrument readings, some of which
showed us a nice breeze this afternoon, and more cubes
of the super pan which is still under construction. Also
received were lander solar array dust monitor images and
end of day stereo images of the rover.

This morning the sojourner rover reposition the APXS, or
Alpha Proton X-Ray Spectrometer, squarely on a somewhat
dust free corner of the rock called "Stimpy". Stimpy is
in the rock garden just to the right of the rock Moe and
Half Dome. We received another close up color image of
Stimpy taken by the rover rear camera which clearly
shows the APXS on the rock Stimpy.

Tomorrow the rover team plans to acquire and send out an
APXS spectrum of Stimpy. Tonight the lander will spend
the night awake for the first time in several days. We
expect to acquire a good deal of over night atmospheric
temperature, pressure, and wind data.

For further information on the Mars Pathfinder Mission,
please call our Mission Status Report line at